OpenStudy (anonymous):

The Doctor has a mixture of 8.5L, x amount of water, and 6L milk , find the equation for milk. the mixture=C, the Water=x and Milk=y C=8.5x+6y The equation for milk is 6y=C-8.5x y=(C-8.5x)/6

OpenStudy (anonymous):

milk = y c= 8.5 + 6y 6y=c-8.5 y= ( c - 8.5)/6

OpenStudy (anonymous):

i forgot to put the x y= (c- 8.5x)/6